OpenStudy (anonymous):

Where does the graph of x2 - x - 2 = y cross the x-axis?

OpenStudy (anonymous):

when y=0, it crosses the x-axis x^2-x-2=0 you can factor this and solve for x

OpenStudy (anonymous):

(x-2)(x+1)=0 x=2, -1 so the points (2,0) and (-1, 0) are x-intercepts